OpenStudy (anonymous):

Can someone help me with this trig question? solve cos theta>= (sqrt3)/2 for 0<=theta<=2pi

OpenStudy (anonymous):

can you draw this or write this in proper mathematical notation?

OpenStudy (anonymous):

Uhm I'm not sure what you mean, this is what my problem asked for. I'll try to rewrite it though. cos θ>/= (squareroot 3)/2 for 0</= θ</= 2pi

OpenStudy (anonymous):

by sqrt 3/2, do you mean this\[\sqrt[3]{2}\] or this \[\sqrt{3/2}\]

OpenStudy (anonymous):

No, its square root of three separately then put over 2

OpenStudy (anonymous):

ok you mean this\[\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

Yeah thats it, I donno how to do the symbols here sorry

OpenStudy (***[isuru]***):

what is the formula that u can use to solve this ? do u know it ?

OpenStudy (anonymous):

No, that's what I don't understand. I used a couple different ones already and they don't work.

OpenStudy (***[isuru]***):

when \[\cos \theta = \ cos\ \alpha \] \[\theta = 2n \pi \pm \alpha\] where n is an integer

OpenStudy (anonymous):

well \[\cos \theta = \frac{ \sqrt{3} }{ 2 }\] for theta = pi/2 or 90 degrees

OpenStudy (***[isuru]***):

that is the equation to solve cos functions so according to that we know in ur case \[\cos \theta = \frac{ \sqrt{3} }{ 2}\] then \[\cos \theta = \cos 30^{0 }\] which means \[\cos \theta = \cos \frac{ \pi }{ 6 }\] so if \[\theta = 2n \pi \pm \alpha\] when \[\cos \theta = \cos \alpha\] then here \[\alpha = \frac{ \sqrt{3} }{ 2 }\] which means \[\theta = 2n \pi \pm \alpha\]\[\theta = 2n \pi \pm \frac{ \sqrt{3} }{ 2}\] now all u have to do is solve it for the given range .Did u got it so far ?

OpenStudy (anonymous):

Okay, I think I get it now. This really helps, thanks ^^

OpenStudy (***[isuru]***):

u r welcome!!